OpenStudy (anonymous):

Solve the system of equations. 9x + y = 2 –4x – y = –17

OpenStudy (anonymous):

Help please?

OpenStudy (cj49):

add them

OpenStudy (cj49):

5x=-15 x=-3

OpenStudy (cj49):

substitute x=-3 in any one of the two equation

OpenStudy (cj49):

9(-3)+y=2 -36+y=2 y=36+2 y=?